Consider the following system of linear equations:

x − y + z = 1

x + y − 2z = 2

2x − z = 3

(a) How many solutions does the system have? Justify your

answer


1
Expert's answer
2021-03-31T17:01:14-0400

Let us find the rank of the extended matrix of the system of equations:

"\left( {\left. {\begin{matrix}
1&{ - 1}&1\\
1&1&{ - 2}\\
2&0&{ - 1}
\end{matrix}} \right|\begin{matrix}
1\\
2\\
3
\end{matrix}} \right)\begin{matrix}
{}\\
{ - I}\\
{ - 2I}
\end{matrix} \sim \left( {\left. {\begin{matrix}
1&{ - 1}&1\\
0&2&{ - 3}\\
0&2&{ - 3}
\end{matrix}} \right|\begin{matrix}
1\\
1\\
1
\end{matrix}} \right)\begin{matrix}
{}\\
{ - II}
\end{matrix} \sim \left( {\left. {\begin{matrix}
1&{ - 1}&1\\
0&2&{ - 3}\\
0&0&0
\end{matrix}} \right|\begin{matrix}
1\\
1\\
0
\end{matrix}} \right)"

Since the ranks of the main and extended matrix are equal to each other, but not equal to the number of equations:

"rank(A) = rank(A|B) = 2 \ne 3" , the system of equations is consistent, but not defined, and therefore has an infinite number of solutions